Express each number as a product of its prime factors: 
 
 5005 

Grade:12

3 Answers

Harshit Singh
3 years ago
Dear student

By Taking the LCM of 5005,
we will get the product of its prime factor.
Hence,
5005 = 5 × 7 × 11 × 13 × 1
= 5 × 7 × 11 × 13

Thanks
KRIPA SHANKAR DWIVEDI
3 years ago
DEAR STUDENTS
First we find the LCM of 5005
  • The product of it’s prime factor is
5005=5×7×11×13×1
5×7×11×13
THANKS
Ram Kushwah
3 years ago
To find the prime factors of 5005 we divide by the prime numbers
2,3,5,7,11,13 etc which ever is possible
 
5005/5=1001
1001/7=143
143/11=13
13/13=1
Thus 5005=5 x 7 x 11 x13